INFRARED INTENSITIES AND SPECTRA OF WATER CLUSTERS

\underline{KIRILL E. KUYANOV}, Department of Chemistry, University of Southern California, Los Angeles, CA 90089; MIKHAIL N. SLIPCHENKO, Department of Chemistry, Iowa State University, Ames, IA 50011; BORIS G. SARTAKOV, General Physics Institute, Russian Academy of Sciences, Moscow, 119991, Russia; ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089.

Water clusters (H_{2}O)_{n} in the range n = 1 – 30 have been studied in helium droplets. Spectral bands of clusters with n = 1 – 6 are observed and infrared intensities of water dimers, trimers and tetramers relative to that of ν_{3} band of H_{2}O molecule have been determined. We found non-additive enhancement of the IR intensity in water clusters, which can be used as a probe for cooperative effects. Gradual transformation of the spectrum from a relatively narrow band of small clusters to a broad spectrum, characteristic for a bulk ice, is observed.